A sample of 500 nursing application included 60 men. Find the 90% confidence interval for the true proportion of men who applied for the nursing program
1
Expert's answer
2020-06-24T17:23:53-0400

since we want 90% confidence interval so "\alpha =0.10" . Since we may be wrong both high or low we need to divide  "\alpha" by 2.

"n=500, x = 60, p = \frac{60}{500} = 0.12, q=1-p = 0.88"


Error term is, "E = Z_{\alpha /2} \sqrt{\frac{pq}{n}} = 1.645 \sqrt{\frac{0.12*0.88}{500}} = 0.02390632722"

so interval is

"0.12 -0.024 < p < 0.12+0.024"

"0.096 < p < 0.144"